• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 23:54:01 on November 22, 1998

    Entry number 6438

    keyword=Beam Time Accounting
    Hall A

    11/22/1998 1600-2400
    Experiment #: E94-010
    Shift Leader: A. Deur
    Shift Workers: R. DiSalvo, A. Gasparian, E. Hughes, H. Ueno

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    35 1 45 35 1 1 50 1 6.75
    Beam Available
    but Not in Use
    25 0 0 25 0 0 10 0 1
    Beam Not Available
    or Unacceptable
    0 0 15 0 0 0 0 0 0.25
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    35 1 45 35 1 1 50 1 6.75
    Planned Config-
    uration Change
    25 0 15 25 0 0 10 0 1.25
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 25 0 0 25 0 0 10 0 1
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 16:00 NMR: 32.3 % Change to perpendicular configration ( @Ep=426 MeV on electron arm and 808 MeV on the hadron) 18:47 Beam went off (MCC have to recycle a magnet) we take this cance to do an NMR (32.4 %) and changed the momentum of electron arm (Ep now at 0.393 GeV) and back to transverse. 22:00 NMR 33.36 % target polarization.

    Data Entry by: A.Deur